OpenStudy (vshiroky):

Find the slope (4,-3) and (-5,-3)

OpenStudy (vshiroky):

I got 0/9 so I'm not sure what that means... is the slope undefined?

OpenStudy (johnweldon1993):

No...it is simply 0....undefined slope would be 9/0 ...we cannot divide by 0...so that is undefined

OpenStudy (♪chibiterasu):

Use the slope formula: \[\frac { y _2 - y _1 }{ x _2 - x _1}\]

OpenStudy (johnweldon1993):

Also remember that would technically be -0/9 right? -3 + 3 0 ------- = ---- -5 - 4 -9 Regardless...0 slope

OpenStudy (vshiroky):

Ok... (-2,3) and (5,7) is 10/7?

OpenStudy (♪chibiterasu):

Yeah, the slope is just 0.

OpenStudy (vshiroky):

I wasn't sure what it meant if the 0 was on top, but I understand it now. Thank you :)

OpenStudy (johnweldon1993):

Not quite... 7 - 3 4 ------ = --- right? 5 - (-2) 7

OpenStudy (vshiroky):

I'm getting ahead of myself... thinking I'm all pro and making mistakes now

OpenStudy (johnweldon1993):

Just remember that slope formula \[\large slope = \frac{y_2 - y_1}{x_2 - x_1}\]